OpenStudy (anonymous):

b/(b^2-25)+5/(b+5)-6/(b) how do I solve this?

OpenStudy (anonymous):

\[b^2 - 25 = (b-5)(b+5)\] \[So, (b^2 + 5b(b-5) - 6(b^2 +5)) / b(b-5)(b+5)\] \[(b^2 + 5b^2 - 25 - 6b^2 -30 ) divb(b-5)(b+5)\] \[-55 \div(b^3-25b)\]

OpenStudy (anonymous):

Do you want to get value for b??

OpenStudy (anonymous):

@TeeTee2292$*

OpenStudy (lgbasallote):

\[\large \Rightarrow \frac{b}{(b+5)(b-5)} + \frac{5}{b+5} - \frac{6}{b}\] \[\large \Rightarrow \frac{b(b) + 5b(b - 5) - 6(b+5)(b-5)}{b(b+5)(b-5)}\] \[\large \Rightarrow \frac{b^2 + 5b^2 - 25b - 6(b^2 - 25)}{b(b^2 - 25)}\] \[\large \Rightarrow \frac{\cancel{6b^2} - 25b \cancel{- 6b^2} + 210}{b(b^2 - 25)}\] \[\large \Rightarrow -\frac{25(b - 6)}{b(b^2 - 25)}\] see where you went wrong @smr_kml ?

OpenStudy (anonymous):

I see Thank you Can you tell me how to write like that?! :)

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

that equation button below the reply box

OpenStudy (anonymous):

@smr_kml you go to eqution in the bottom left of the reply box

OpenStudy (lgbasallote):

for fractions type this in the equation editor \frac{numerator}{denominator} this results in \(\frac{numerator}{denominator}\)

OpenStudy (anonymous):

ok thank you @lgbasallote and @parthdesai